20 February 2014 – Thursday

LDAP:

  • One issue in last night’s update.  A former student.  The dance was done.
  • Also discovered a left over from Valentine’s day.  That was the new faculty that has a singleton match with a Lecturer, but has a different last name.  The Registrar’s office has never finished confirming that one.  Followed up with the Registrar’s office and was told that the match and need to change the name had been confirmed.  So, I merged the two records in LDAP.
  • writing possible VPN group munging code to handle appearing and disappearing accounts.  Ran ok (with no SOR changes) on the test system.  Going to leave it running there and we’ll see how it does.

Backups:

  • One group is still running that should not be.  The single client hanging out, says it is “state started”.  Asking the administrator to check up on it and see if it is really started, and if they can’t find any indication that it is, would they restart the networker service on that windows system.
  • Clones:  CoM/IS has 51 SSIDS, 18.3GB left (working on 18, 8.3 now).  Disk has 91 SSIDS, 16,244 GB (working on 1, 902.1 now).
  • The shred of the first of the two 35TB LUNs from the Nexsan SATABeasts is 5%  (1.8TB) done on its first pass writing random junk to all the disks after a day and a half…

Tags:

Leave a Reply

You must be logged in to post a comment.